Value-Stream Average Costing, ABC Costs as Benchmarks

A value stream has three activities and two products. The units produced and shipped per week are 50 of the deluxe model (Model A) and 150 of the basic model (Model B). The resource consumption patterns are shown as follows:


Model A

Model B
Costs of Value-
Stream Activities
Cell manufacturing 2,395 min. 7,185 min. $ 57,480
Engineering 70 hrs. 280 hrs. 30,800
Testing 95 hrs. 285 hrs. 29,640
   Total $ 117,920

Required:

1. Calculate the ABC product cost for Models A and B. If required, round your answers to the nearest cent.

Product Cost Per Unit
Model A $ per unit
Model B $ per unit

2a. Calculate the value-stream average product cost. If required, round your answer to the nearest cent.
$ per unit

2b. Assuming reasonable stability in the consumption patterns of the products and product mix, assess how well the products are grouped based on similarity.
Answer #1

Ans:

1)ABC product cost for model A & B

Total cost = (Cost of value stream activities / Total activity)* relevant model activity

Product cost per unit = total cost / total units

Model A

Total cost =

((57,480/(2395+7185))*2395+((30,800/(70+280))*70)+(29,640/(95+285))*95))=(14370+6160+7410)

= 27,940

Product cost per unit = 27,940 / 50 = $ 558.8

Model B

Total cost =

((57,480/(2395+7185))*7185+((30,800/(70+280))*280)+(29,640/(95+285))*285))

= (43110+ 24640 + 22230)

= 89,980

Product cost per unit = 89,980 / 150 = $599.87

2) a. Average product cost = total cost for product A and B / Total units

  Average product cost = (27940+ 89,980) / (50 + 150)

= 117920 / 200

= $589.6

The two value –Stream Products are similar